# SeeGround: See and Ground for Zero-Shot Open-Vocabulary 3D Visual Grounding

### v1

@@ Abstract
3D Visual Grounding (3DVG) aims to locate objects in 3D scenes based on textual descriptions, which is essential for applications like augmented reality and robotics. Traditional 3DVG approaches rely on annotated 3D datasets and predefined object categories, limiting scalability and adaptability. To overcome these limitations, we introduce SeeGround, a zero-shot 3DVG framework leveraging 2D Vision-Language Models (VLMs) trained on large-scale 2D data. We propose to represent 3D scenes as a hybrid of query-aligned rendered images and spatially enriched text descriptions, bridging the gap between 3D data and 2D-VLMs input formats. We propose two modules: the Perspective Adaptation Module, which dynamically selects viewpoints for query-relevant image rendering, and the Fusion Alignment Module, which integrates 2D images with 3D spatial descriptions to enhance object localization. Extensive experiments on ScanRefer and Nr3D demonstrate that our approach outperforms existing zero-shot methods by large margins. Notably, we exceed weakly supervised methods and rival some fully supervised ones, outperforming previous SOTA by 7.7% on ScanRefer and 7.1% on Nr3D, showcasing its effectiveness in complex 3DVG tasks.
@@ 1 Introduction
3D Visual Grounding (3DVG) aims to locate specific objects within a 3D scene based on given textual descriptions, playing a crucial role in applications such as augmented reality [4], vision-language navigation [6], and robotic perception [5, 22, 21]. Effective solutions require both textual comprehension and spatial reasoning capabilities within diverse and complex 3D environments.
@@
To address these limitations, we propose SeeGround, leveraging 2D Vision-Language Models (VLMs) [28, 37, 12] for flexible 3DVG. Trained on extensive 2D data, 2D-VLMs offer open-vocabulary understanding ability, performing well in tasks like image captioning [19] and visual question answering [52]. This capability provides insight for zero-shot 3DVG. Considering that 2D-VLMs cannot process 3D data directly, we introduce a cross-modal alignment representation that enables 2D-VLMs to interpret 3D scenes. This approach combines 2D rendered images with spatially enriched text descriptions, aligning visual and spatial information so that the 2D-VLMs can comprehend the 3D structure and relationships within the scene, thereby achieving more accurate object grounding in the diverse and complex 3D scenes.
Specifically, we propose to represent 3D scenes as a combination of “2D rendered images” and “3D spatial descriptions”. Specifically, images are rendered using query-driven dynamic viewpoints, simulating relevant observation angles, and capturing object details and spatial context. This approach avoids redundancy in multi-view methods and limitations of bird’s-eye views, which lack height and orientation details. The 3D spatial descriptions from pre-saved object detection provide accurate 3D positional information, enhancing the 2D-VLMs’ understanding of object relationships within the scene.
However, when textual descriptions and images are processed separately by 2D-VLMs, the model cannot associate 3D spatial information from text to the object in the 2D images. For example, in scenes with multiple similar objects (such as several chairs), it can be challenging for the model to identify which chair corresponds to the specified one on the image. To address this, we introduce a new visual prompting technique that explicitly marks key objects within images, establishing clear correspondences between 2D images and 3D spatial descriptions. Visual prompting not only enhances the fusion of visual and spatial information but also guides the model’s attention to target areas, reducing potential interference from irrelevant information and improving the efficiency and accuracy of localization in multi-object 3D scenes.
@@
  • We introduce SeeGround, a training-free solution for zero-shot 3DVG. It converts 3D scenes into 2D-VLMs compatible format and combines 2D rendered images with 3D spatial descriptions, leveraging the open-vocabulary capabilities of 2D-VLMs to achieve zero-shot 3DVG without reliance on 3D-specific training data.
@@ 2 Related work
3D Visual Grounding. Supervised 3DVG methods, e.g., ScanRefer [4] and ReferIt3D [1], achieve object localization by aligning 3D scenes with natural language descriptions. 3DVG-Transformer [51] further refined localization through attention mechanisms. ViewRefer [11] extends input text with large language models for multi-view semantic capture, while MVT [15] and LAR [2] leverage spatial context from multi-perspective views. SAT [46] introduces 2D semantic-assisted training to better align 2D-3D features, and BUTD-DETR [17] combines bottom-up and top-down detection with transformers. ConcreteNet [36] frames grounding as 3D instance segmentation, and WS-3DVG [41] uses limited annotations in a coarse-to-fine matching strategy. While supervised methods excel on benchmarks, they require extensive annotations, limiting scalability. Recent zero-shot methods, such as LLM-Grounder [44] and ZSVG3D [49], are annotation-free and enhance adaptability. However, text-driven approaches often miss critical visual details impacting precise localization.
@@
3D Open-Vocabulary Understanding. Recent advances in 3D scene understanding enable open-vocabulary capabilities for tasks like segmentation and retrieval through 2D-3D alignments [5]. OpenScene [30] achieves open-vocabulary segmentation by projecting 2D pixel-wise features onto 3D scenes. LeRF [20] integrates multi-scale CLIP features into a neural radiance field. OVIR-3D [27] merges multi-view 2D region proposals into 3D space for open-vocabulary instance retrieval. Agent3D-Zero [50] uses vision-language models across multiple perspectives for 3D question answering and segmentation. RegionPLC [45] generates 3D-text pairs with 2D captions. OpenMask3D [35] uses aligned images to propose masks for open-vocabulary instance segmentation. OpenIns3D [16] enhances segmentation across varied 3D scenes through synthetic data. SAI3D [47] employs Semantic-SAM to acquire 2D masks, connecting them to 3D regions with graph-based merging.
@@
MLLMs Models for 3D Perception. Recent progress in MLLMs has enhanced 3D understanding by applying 2D techniques to 3D contexts [26, 25, 43]. Scene-LLM [10] expands MLLMs capabilities by supporting 3D dense captioning and semantic segmentation. Uni3DL [23] introduces a unified framework combining 3D understanding with language comprehension, while 3D-ViSTA [53] leverages transformers to align 3D visual data with text inputs, advancing dual-modality comprehension. ConceptFusion [18] integrates 3D object instances with conceptual knowledge from language, reinforcing 3D semantic understanding and enabling reasoning. SceneVerse [19] introduces a language-annotated dataset of 3D environments, helping MLLMs learn spatial relationships. In addition, RLHF-V [33] enables agents to perform 3D tasks from natural language commands, supporting interactive tasks such as ac and task planning. These models highlight MLLMs’ adaptability in enhancing 3D perception, reasoning, and spatial understanding. Our approach builds on these advancements by providing a zero-shot 3D grounding model that enhances adaptability, enhances robustness, and offers broader viewpoint coverage for complex 3D tasks.
@@
Figure 2: Overview of the SeeGround framework. We first use a 2D-VLM to interpret the query, identifying both the target object (e.g., “laptop”) and a context-providing anchor (e.g., “chair with floral pattern”). A dynamic viewpoint is then selected based on the anchor’s position, enabling the capture of a 2D rendered image that aligns with the query’s spatial requirements. Using the Object Lookup Table (OLT), we retrieve the 3D bounding boxes of relevant objects, project them onto the 2D image, and apply visual prompts to mark visible objects, filtering out occlusions. The image with prompts, along with the spatial descriptions and query, are then input into the 2D-VLM for precise localization of the target object. Finally, the 2D-VLM outputs the target object’s ID, and we retrieve its 3D bounding box from the OLT to provide the final, accurate 3D position in the scene.
@@ 3 Methodology
3.1 Overview
[…]
$\text{bbox}=\mathbf{3DVG}(\mathcal{S},\mathcal{Q}).$ (1)
In this work, we propose a novel method for 3DVG that integrates 2D-VLM with spatially enriched 3D scene representations. Traditional 3D scene models are not directly compatible with the input format required by 2D-VLM, which have shown great promise in scene understanding tasks like image captioning and visual question answering [19, 37]. To bridge this gap, we introduce a hybrid representation: it utilizes 2D rendered images that can be easily processed by 2D-VLM, while also incorporating text-based 3D spatial descriptions. This representation allows our framework to align the rich visual features from 2D renderings with the spatial context from 3D scene descriptions. By doing so, we facilitate effective multimodal information alignment and ensure that the 2D-VLM can understand and reason about objects in complex 3D environments without the need for additional 3D-specific training.
We first establish a multimodal 3D representation that is compatible with the 2D-VLM input format in Sec. 3.2, comprising a perspective-aligned rendered image and a spatial description in the text. For each query, the Perspective Adaptation Module (Sec. 3.3) generates a 2D rendered image from a perspective that aligns with the query, capturing relevant objects and spatial relationships. The spatial description, stored in the Object Lookup Table (OLT), includes the 3D bounding boxes and semantic labels of these objects. Further, the Fusion Alignment Module (Sec. 3.4) integrates the rendered image with the 3D spatial descriptions, creating an aligned multimodal representation. This alignment allows the 2D-VLM to process the query, the aligned image, and the text description together, enabling precise localization of the target object and outputting the target object. The overall framework is illustrated in Fig. 2.
@@ 3.2 Multimodal 3D Representation
Our method proposes to leverage 2D-VLM, which are trained on large-scale text-image datasets to capture rich prior knowledge. This enables the model to interpret novel objects and scenes, facilitating open-set understanding. However, prior 3D scene representations – such as point clouds [30, 13], voxels [24], and implicit representations [20] – are not directly compatible with the input format required by 2D-VLM. To bridge this gap, a 3D scene representation that aligns with 2D-VLM input is necessary. To tackle this problem, in this work, we propose a hybrid representation that combines “2D rendered images” and “text-based 3D spatial descriptions”.
Text-based 3D Spatial Descriptions. The process begins with detecting all objects in the 3D scene. Using an open-vocabulary 3D detection framework, we identify each object’s 3d bounding boxes bbox (positions and sizes), and semantic labels sem. This can be formulated as: $(\text{bbox},\text{sem})_{i=1}^{N}=\mathbf{OVDet}(\mathcal{S})$. This information is then converted into a text format compatible with the 2D-VLM input, allowing an accurate spatial and semantic description of the scene. Since a single scene can correspond to multiple queries, object detection is performed only once per scene, and all $N$ detected objects are stored in the OLT for efficient processing of subsequent queries. Additionally, the OLT enables the model to retrieve spatial information efficiently, avoiding complex spatial relationship calculations in later steps. It is formally defined as:
$\mathcal{OLT}=\left\{\left(\text{bbox},\text{sem}\right)\right\}_{i=1}^{N}.$ (2)
@@
$\left(\mathcal{I},\mathcal{T}\right)=\mathcal{F}\left(\mathcal{S},\mathcal{Q},\mathcal{OLT}\right),$ (3)
where $\mathcal{F}$ takes the 3D scene, the query $\mathcal{Q}$ and the $\mathcal{OLT}$ as inputs, to generate a 2D rendered image $\mathcal{I}$ and a text-based spatial description $\mathcal{T}$.
The text-based spatial descriptions provide each object’s accurate 3D position, size, and semantic label, assisting the 2D-VLM in understanding basic spatial relationships between objects. The rendered images offer a 2D perspective of the 3D scene, allowing the model to capture visual features such as color, shape, and texture, along with relative spatial positions. By combining them, the 2D-VLM gains a comprehensive understanding of both the visual details and true 3D spatial information within the scene.
@@ 3.3 Perspective Adaptation Module
Dynamic Perspective Selection. This process is guided by contextual information in the text instruction $\mathcal{Q}$. The 2D-VLM identifies the anchor object $\mathcal{A}$ and candidate target objects $\mathcal{O}^{(C)}$ by analyzing relationships in $\mathcal{Q}$. We provide example prompts $\mathcal{E}^{(E)}$ to help the model understand these relationships. This process is formalized as:
$\left(\mathcal{A},\mathcal{O}^{(C)}\right)=\text{VLM}\left(\mathcal{Q},\mathcal{E}^{(E)}\right).$ (4)
Based on the identified anchor $\mathcal{A}$, a rendering viewpoint is selected to capture the scene from a perspective aligned with the query. The viewpoint is initially positioned at the center of the 3D scene, focusing on the chosen anchor $\mathcal{A}$. The camera then moves backward and upward, away from the anchor, to gradually cover a broader view of the scene. In the case where there is no anchor object identified during the analysis (e.g., the query describes multiple similar objects), a placeholder anchor is introduced, using the target object as the substitute. If multiple targets are present, the center point of these targets is used as the placeholder anchor. The subsequent perspective selection steps then proceed as described earlier.
Query-Aligned Image Rendering. Once the perspective is determined, a virtual camera is placed at this position, and the function look_at_view_transform (detailed in the Supplementary Material) generates the rotation matrix $\mathcal{R}_{c}$ and translation vector $\mathcal{T}_{c}$ based on the virtual camera’s position and orientation relative to the anchor $\mathcal{A}$. Using these matrices, the scene point cloud is rendered into a 2D image aligned with the query’s description, represented by:
$\mathcal{I}=\mathbf{Render}(\mathcal{S},\mathcal{R}_{c},\mathcal{T}_{c}).$ (5)
This approach enables the scene to be observed from a query-consistent viewpoint, providing clear visual details that avoid potential misinterpretations by the 2D-VLM. Additionally, as shown in Fig. 3 (e), filtering irrelevant information enhances localization accuracy by reducing interpretive confusion within the model.
@@
Table 1: Evaluations of 3DVG on ScanRefer [4] validation set. Results are reported for “Unique” (scenes with a single target object) and “Multiple” (scenes with distractors of the same class) subsets, along with overall performance.
@@
SeeGround Ours Zero-Shot Qwen2-VL-72b [37] $\mathbf{75.7}$ $\mathbf{68.9}$ $\mathbf{34.0}$ $\mathbf{30.0}$ $\mathbf{44.1}$ $\mathbf{39.4}$

### v2

@@ Abstract
3D Visual Grounding (3DVG) aims to locate objects in 3D scenes based on textual descriptions, essential for applications like augmented reality and robotics. Traditional 3DVG approaches rely on annotated 3D datasets and predefined object categories, limiting scalability and adaptability. To overcome these limitations, we introduce SeeGround, a zero-shot 3DVG framework leveraging 2D Vision-Language Models (VLMs) trained on large-scale 2D data. SeeGround represents 3D scenes as a hybrid of query-aligned rendered images and spatially enriched text descriptions, bridging the gap between 3D data and 2D-VLMs input formats. We propose two modules: the Perspective Adaptation Module, which dynamically selects viewpoints for query-relevant image rendering, and the Fusion Alignment Module, which integrates 2D images with 3D spatial descriptions to enhance object localization. Extensive experiments on ScanRefer and Nr3D demonstrate that our approach outperforms existing zero-shot methods by large margins. Notably, we exceed weakly supervised methods and rival some fully supervised ones, outperforming previous SOTA by 7.7% on ScanRefer and 7.1% on Nr3D, showcasing its effectiveness in complex 3DVG tasks. Project website (with demo and code): https://seeground.github.io.
@@ 1 Introduction
3D Visual Grounding (3DVG) aims to locate specific objects within a 3D scene based on given textual descriptions, playing a crucial role in applications such as augmented reality [5, 36, 35, 37, 57, 39], vision-language navigation [8, 20, 13], and robotic perception [6, 27, 26, 28, 29, 72, 50, 30, 73, 17]. Effective solutions require both textual comprehension and spatial reasoning capabilities within complex 3D environments.
@@
To address these limitations, we propose SeeGround, leveraging 2D Vision-Language Models (VLMs) [41, 52, 15] for flexible 3DVG. Trained on extensive 2D data, 2D-VLMs offer open-vocabulary understanding ability, performing well in tasks like image captioning [24] and visual question answering [69]. This capability provides insight for zero-shot 3DVG. Considering that 2D-VLMs cannot process 3D data directly, we introduce a cross-modal alignment representation that enables 2D-VLMs to interpret 3D scenes. This approach combines 2D rendered images with spatially enriched text descriptions, aligning visual and spatial information so that the 2D-VLMs can comprehend the 3D structure and relationships within the scene, thereby achieving more accurate object grounding.
Specifically, we propose to represent 3D scenes as a combination of “2D rendered images” and “3D spatial descriptions”. The images are rendered using query-driven dynamic viewpoints, simulating relevant observation angles, and capturing object details and spatial context. This approach avoids redundancy in multi-view methods and limitations of bird’s-eye views, which lack height and orientation details. The 3D spatial descriptions from pre-saved object detection provide accurate 3D positions, enhancing the VLMs’ understanding of object relationships.
However, when textual descriptions and images are processed separately by 2D-VLMs, the model cannot associate 3D spatial information from text to the object in the 2D images. For example, in scenes with multiple similar objects (e.g., several chairs), it can be challenging for the model to identify which chair corresponds to the specified one on the image. To address this, we introduce a new visual prompting technique that explicitly marks key objects within images, establishing clear correspondences between 2D images and 3D spatial descriptions. Visual prompting not only enhances the fusion of visual and spatial information but also guides the model’s attention to target areas, reducing potential interference from irrelevant information and improving the localization accuracy in multi-object 3D scenes.
@@
  • We introduce SeeGround, a training-free solution for zero-shot 3DVG. It converts 3D scenes into 2D-VLMs compatible format and combines 2D rendered images with 3D spatial descriptions, leveraging the open-vocabulary capabilities of 2D-VLMs to achieve zero-shot 3DVG without depending on 3D-specific training data.
@@ 2 Related work
3D Visual Grounding. Supervised 3DVG methods, e.g., ScanRefer [5] and ReferIt3D [1], achieve object localization by aligning 3D scenes with natural language descriptions. 3DVG-Transformer [68] further refined localization through attention mechanisms. More recent work explores enhanced multimodal integration: ViewRefer [14] extends input texts with large language models for multi-view semantic capture. MVT [19] and LAR [2] leverage spatial context from multi-perspective views. SAT [63] introduces 2D semantic-assisted training to better align 2D-3D features. BUTD-DETR [22] combines bottom-up and top-down detection with transformers. ConcreteNet [51] frames grounding as 3D instance segmentation. WS-3DVG [56] uses limited annotations in a coarse-to-fine matching strategy. PQ3D [71] proposes a unified framework for multiple 3D-VL tasks. While supervised methods excel on benchmarks, they require extensive annotations, limiting scalability. Recent zero-shot methods, such as LLM-Grounder [61] and ZSVG3D [66], are annotation-free and enhance adaptability. However, text-driven approaches often miss critical visual details that could impact precise 3D localization.
@@
3D Open-Vocabulary Understanding. Recent advances in 3D scene understanding enable open-vocabulary capabilities for tasks like segmentation and retrieval through 2D-3D alignments [6]. OpenScene [43] achieves open-vocabulary segmentation by projecting 2D pixel-wise features onto 3D scenes. LeRF [25] integrates multi-scale CLIP features into a neural radiance field. OVIR-3D [38] merges multi-view 2D region proposals into 3D space for open-vocabulary instance retrieval. Agent3D-Zero [67] uses vision-language models across multiple perspectives for 3D question answering and segmentation. RegionPLC [62] generates 3D-text pairs with 2D captions. OpenMask3D [49] uses aligned images to propose masks for open-vocabulary instance segmentation. OpenIns3D [21] enhances segmentation across varied 3D scenes through synthetic data. SAI3D [64] employs Semantic-SAM to acquire 2D masks, connecting them to 3D regions with graph-based merging. These approaches demonstrate the value of 2D-to-3D alignment for open-vocabulary understanding in 3D scenes.
@@
MLLMs Models for 3D Perception. Recent progress in MLLMs has enhanced 3D understanding by applying 2D techniques to 3D contexts [34, 33, 60]. Scene-LLM [12] expands MLLMs’ capabilities by supporting 3D dense captioning and semantic segmentation. Uni3DL [31] introduces a unified framework combining 3D understanding with language comprehension, while 3D-ViSTA [70] leverages transformers to align 3D visual data with text inputs, advancing dual-modality comprehension. ConceptFusion [23] integrates 3D object instances with conceptual knowledge from language, reinforcing 3D semantic understanding and enabling reasoning. Glover [40] handles 3D manipulation task. SceneVerse [24] introduces a language-annotated dataset of 3D environments, helping MLLMs to learn spatial relationships. In addition, RLHF-V [47] enables agents to perform 3D tasks from natural language commands, supporting interactive tasks such as ac and task planning. These models highlight MLLMs’ adaptability in enhancing 3D perception, reasoning, and spatial understanding. Our approach builds on these advancements by offering a zero-shot 3D grounding model with improved adaptability, robustness, and broader viewpoint coverage for complex 3D tasks.
@@
Figure 2: Overview of the SeeGround framework. We first use a 2D-VLM to interpret the query, identifying both the target object (e.g., “laptop”) and a context-providing anchor (e.g., “chair with a floral pattern”). A dynamic viewpoint is then selected based on the anchor’s position, enabling the capture of a 2D rendered image that aligns with the query’s spatial requirements. Using the Object Lookup Table ($\mathcal{OLT}$), we retrieve the 3D bounding boxes of relevant objects, project them onto the 2D image, and apply visual prompts to mark visible objects, filtering out occlusions. The image with prompts, along with the spatial descriptions and query, is then input into the 2D-VLM for precise localization of the target object. Finally, the 2D-VLM outputs the target object’s ID, which is then used to retrieve its 3D bounding box from the $\mathcal{OLT}$, providing the final, accurate 3D position in the scene.
@@ 3 Methodology
Overview. The task of 3DVG aims to precisely locate a target object within a 3D scene $\mathcal{S}$, based on a textual description $\mathsf{Q}$. The goal is to output a directed 3D bounding box $(\mathbf{bbox})$ of object $o$ that identifies the target object’s location and dimensions. Formally, this process can be expressed as:
$\mathbf{bbox}=\mathbf{3DVG}(\mathcal{S},\mathsf{Q}).$ (1)
In this work, we propose a novel method for 3DVG that integrates 2D-VLM with spatially enriched 3D scene representations. Traditional 3D scene models are not directly compatible with the input format required by 2D-VLM, which has shown great promise in scene understanding tasks like image captioning and visual question answering [24, 52]. To bridge this gap, we introduce a hybrid representation: it utilizes 2D rendered images that are easily processed by 2D-VLM, while also incorporating text-based 3D spatial descriptions. This representation allows our framework to align the rich visual features from 2D renderings with the spatial context from 3D scene descriptions. By doing so, we facilitate effective multimodal information alignment and ensure that the 2D-VLM can understand and reason about objects in complex 3D environments without the need for additional 3D-specific training.
We first establish a multimodal 3D representation that is compatible with the 2D-VLM input format in Sec. 3.1, comprising a perspective-aligned rendered image and a spatial description in the text. For each query, the Perspective Adaptation Module (Sec. 3.2) generates a 2D rendered image from a perspective that aligns with the query, capturing relevant objects and spatial relationships. The spatial description, stored in the Object Lookup Table ($\mathcal{OLT}$), includes the 3D bounding boxes and semantic labels of these objects. Further, the Fusion Alignment Module (Sec. 3.3) integrates the rendered image with the 3D spatial descriptions, creating an aligned multimodal representation. This alignment allows the 2D-VLM to process the query, the aligned image, and the text description together, enabling precise localization and retrieval of the target object. The overall framework is illustrated in Fig. 2.
@@ 3.1 Multimodal 3D Representation
Our method proposes to leverage 2D-VLM, which is trained on large-scale text-image datasets to capture rich prior knowledge. This enables the model to interpret novel objects and scenes, facilitating open-set understanding. However, prior 3D scene representations – such as point clouds [43, 16], voxels [32], and implicit representations [25] – are not directly compatible with the input format required by 2D-VLM. To bridge this gap, a 3D scene representation that aligns with 2D-VLM input is necessary. To tackle this problem, in this work, we propose a hybrid representation that combines “2D rendered images” and “text-based 3D spatial descriptions”.
Text-based 3D Spatial Descriptions. The process begins with detecting all objects in the 3D scene. Using an open-vocabulary 3D detection framework, we identify each object’s 3d bounding boxes $\mathbf{bbox}$ (positions and dimensions), and semantic labels $\mathbf{sem}$. This can be formulated as: $(\mathbf{bbox},\mathbf{sem})_{i=1}^{N}=\mathbf{OVDet}(\mathcal{S})$. This information is then converted into a text format compatible with the 2D-VLM input, allowing an accurate spatial and semantic description of the scene. Since a single scene may correspond to multiple queries, object detection is performed only once per scene, and all $N$ detected objects are stored in the $\mathcal{OLT}$ for efficient processing of subsequent queries. Additionally, the $\mathcal{OLT}$ enables the model to retrieve spatial information efficiently, avoiding complex spatial relationship calculations in later steps. It is formally defined as follows:
$\mathcal{OLT}=\left\{\left(\mathbf{bbox},\mathbf{sem}\right)\right\}_{i=1}^{N}.$ (2)
@@
$\left(\mathbf{I},\mathcal{T}\right)=\mathbf{F}\left(\mathcal{S},\mathsf{Q},\mathcal{OLT}\right),$ (3)
where $\mathbf{F}$ takes the 3D scene $\mathcal{S}$, the query $\mathsf{Q}$ and the object lookup table $\mathcal{OLT}$ as inputs, to generate a 2D rendered image $\mathbf{I}$ and a text-based spatial description $\mathcal{T}$.
The text-based spatial descriptions provide each object’s accurate 3D location, dimension, and semantic label, assisting the 2D-VLM in understanding basic spatial relationships between objects. The rendered images offer a 2D perspective of the 3D scene, allowing the model to capture visual features such as color, shape, texture, and relative spatial positions. By combining them, the 2D-VLM gains a comprehensive understanding of both the visual details and true 3D spatial information within the scene.
@@ 3.2 Perspective Adaptation Module
Dynamic Perspective Selection. This process is guided by contextual information in the text instruction $\mathsf{Q}$. The 2D-VLM identifies the anchor object $\boldsymbol{A}$ and candidate target objects $\mathcal{O}^{(C)}$ by analyzing relationships in $\mathsf{Q}$. We provide example prompts $\mathcal{E}^{(E)}$ to help the model understand these relationships (see Supp. for prompt construction details). This process is formalized as:
$\left(\boldsymbol{A},\mathcal{O}^{(C)}\right)=\operatorname{VLM}\left(\mathsf{Q},\mathcal{E}^{(E)}\right).$ (4)
Based on the identified anchor $\boldsymbol{A}$, a rendering viewpoint is selected to capture the scene from a perspective aligned with the query. The viewpoint is initially positioned at the center of the 3D scene, focusing on the chosen anchor $\boldsymbol{A}$. The camera then moves backward and upward, away from the anchor, to gradually cover a broader view of the scene. In the case where there is no anchor object identified during the analysis, e.g., the query describes multiple similar objects), a placeholder anchor is introduced, with the target object serving as the substitute. If multiple targets are present, the center point of these targets is used as the placeholder anchor. The subsequent perspective selection steps then proceed as described earlier.
Query-Aligned Image Rendering. Once the perspective is determined, a virtual camera is placed at this position, and the function $\operatorname{look\_at\_view\_transform}$ (see Supp. for details) generates the rotation matrix $\mathbf{R}_{c}$ and translation vector $\mathbf{T}_{c}$ based on the virtual camera’s position and orientation relative to the anchor $\boldsymbol{A}$. With these matrices, the scene point cloud is projected and rendered into a 2D image aligned with the query description, represented by:
$\mathbf{I}=\operatorname{Render}(\mathcal{S},\mathbf{R}_{c},\mathbf{T}_{c}).$ (5)
This approach enables the scene to be observed from a query-consistent viewpoint, providing clear visual details that avoid potential misinterpretations by the 2D-VLM. Additionally, as shown in Fig. 3 (e), filtering out irrelevant information enhances localization accuracy by reducing interpretive confusion within the model.
@@
Table 1: Evaluations of 3DVG on ScanRefer [5] validation set. Results are reported for “Unique” (scenes with a single target object) and “Multiple” (scenes with distractors of the same class) subsets, along with overall performance. * indicates results on selected 250 samples.
@@
VLM-Grounder* [59] CoRL’24 Zero-Shot GPT-4V [41] $66.0$ $29.8$ $\mathbf{48.3}$ $\mathbf{33.5}$ $\mathbf{51.6}$ $32.8$
SeeGround Ours Zero-Shot Qwen2-VL-72b [52] $\mathbf{75.7}$ $\mathbf{68.9}$ $34.0$ $30.0$ $44.1$ $\mathbf{39.4}$